Circular No. 5809

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993Q IN ANONYMOUS GALAXY R. H. McNaught, Siding Spring, reports his discovery of a supernova

(mag 19) on a U.K. Schmidt Telescope R plate taken on May 28.8 UT by M. Hartley. The supernova is located at α = 20^{h}32^{m}24^{s}.61, δ = −42^{o}57′57′′.7 (equinox 1950.0); the galaxy’s nucleus is located at (end figures) 23^{s}.76, 54′′.9, and a star of mag 18.5 is at 22^{s}.76, 43′′.4. The positions are probably good to ± 1′′.5, as the object is situated in the corner of the plate. The supernova is not present on SERC J or ESO R surveys.

M. Della Valle, European Southern Observatory, reports: “The analysis of a spectrogram (range 400-840 nm, resolution ∼ 2 nm), obtained at La Silla with the New Technology Telescope on May 30.3 UT (under nonphotometric conditions, reveals this object to be a type-Ia supernova 2-3 weeks past maximum. The expansion velocity deduced from the minimum of the Si II 635-nm absorption is ∼ 9000 km/s.”

SUPERNOVA 1993J IN NGC 3031 S. Benetti and R. Barbon, Asiago Astrophysical Observatory, commu-

nicate: “A CCD grating spectrum of SN 1993J at 2.3-nm resolution, taken on May 28.89 UT with the 1.82-m Cima Ekar telescope, shows a complex structure with conspicuous emissions at 864.8 (Ca II IR), 777.3 (O I), 728.2 (Ca II]), 704.2 ([Fe II]?), 663.7 and 647.3 (structured Hα), 626.0 ([O I]), 586.8 (Na I with a strong P-Cyg profile), and 553.4 nm ([O I]?, [Fe II]?). Many fainter emissions are also seen in the blue region up to H+K Ca II absorption at 383.7 nm. Also Hβ seems to show the same Hα structure, although blends with the Fe II lines cannot be excluded. The appearance of forbidden lines of [O I] and Ca II] suggests that the supernova is already entering the nebular phase.”

Further photoelectric photometry by H. Mikuž, B. Dintinjana, and T. Zwitter, Ljubljana, Slovenia (cf. IAUC 5796): Apr. 8.90, B = 12.08 ± 0.01, V = 11.64 ± 0.03, R = 11.12 ± 0.01, I = 11.02 ± 0.03; 17.90, –, –, 10.46 ± 0.02, 10.37 ± 0.02; 17.97, 11.37 ± 0.02, 10.80 ± 0.01, 10.44 ± 0.03, 10.41 ± 0.03; 20.93, 11.74 ± 0.01, 10.83 ± 0.01, 10.56 ± 0.03, 10.41 ± 0.03; 22.86, 11.82 ± 0.05, 11.04 ± 0.02, 10.69 ± 0.03, 10.51 ± 0.03; 23.91, 12.01 ± 0.02, 11.13 ± 0.01, 10.65 ± 0.02, 10.34 ± 0.04; 26.88, 12.46 ± 0.05, 11.43 ± 0.02, 10.92 ± 0.02, –; May 17.89, 13.47 ± 0.04, 12.37 ± 0.02, 11.77 ± 0.01, 11.29 ± 0.03; 22.94, 13.67 ± 0.05, 12.43 ± 0.02, 11.83 ± 0.02, 11.47 ± 0.03.

1993 May 31 Daniel W. E. Green